OpenStudy (anonymous):

HELP |2x − 7| ≤ 0.2 Whats the interval notation set? I got 2x-7 ≤ 0.2 and 2x - 7 ≤ -0.2 x ≤ 3.6 and x≤3.4 so... x≤3.4?

OpenStudy (anonymous):

I got 2x-7 ≤ 0.2 and 2x - 7 ≤ -0.2 x ≤ 3.6 and x≤3.4 so... x≤3.4?

OpenStudy (anonymous):

beginning is wrong

OpenStudy (anonymous):

\[|2x-7|\leq 0.2\iff -0.2\leq 2x-7\leq 0.2\] add 7 divide by 2

OpenStudy (anonymous):

Finally someone answered. Thanks. so that would be... 2.5<(equal) x <(equal) 3.6 right?

OpenStudy (anonymous):

3.4**<(equal)...

OpenStudy (anonymous):

Is this going to be the same for all questions like this?